Sol Summary – April 26th

Crew 280 Sol Summary Report 26-04-2023

Sol: 10

Summary title: Harvesting Martian results

Author’s name: Mariona Badenas-Agustí, Commander and Astronomer

Mission status: Nominal

Sol activity summary:

The Hypatia I crew started their day with enthusiasm, eager to embark on their longest EVA yet. After gathering around the table for breakfast, Carla Conejo (Executive Officer & Biologist), Núria Jar (Journalist), Laia Ribas (GreenHab Officer), and Mariona Badenas-Agusti (Crew Commander & Astronomer) prepared themselves for their EVA. Today, they drove towards Brahe Highway 1572, stopping just before reaching the Martian moons. In their final destination, they flew a small drone and collected sand samples for Carla’s and Laia’s project, respectively.

The full crew gathered at lunchtime, when Cesca Cufí (Engineer) and Neus Sabaté (Mission Specialist) prepared a refreshing rice salad with beans, dehydrated chicken (or mushrooms for vegetarians), and fresh chives from the GreenHab. After lunch, the crew resumed their work on their experiments in the Science Dome, RAM, and the Hab. As always, it has been a productive day on Mars.

With only two more SOLs left before their return to Earth, the Hypatia I crew can’t help but feel a sense of nostalgia about their time on Mars and how quickly their mission is going by. Nevertheless, they remain highly motivated to complete their projects and share their findings with those who have followed their journey, both back home and beyond.

Look ahead plan: Tomorrow will be characterized by two EVAs dedicated to a navigation competition between the crew members (See EVA Request).

Anomalies in work: None.

Weather: Sunny. Low 38F/High 66F.

Crew physical status: Nominal.

EVA: There was an EVA to Cow Dung Road 0110 (see EVA Report).

Reports to be filed: Sol Summary, EVA request, EVA Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.

Support Requested: None.

Operations Report – April 26th

Crew 280 Operations Report 26-04-2023
SOL: 10
Name of person filing report: Neus Sabaté (Mission Specialist)
Non-nominal systems:
Robotic Observatory
Notes on non-nominal systems: Replacement parts of Robotic observatory are scheduled to be installed during crew 261 rotation.
Spirit rover used: yes
Hours: 219.8
Beginning charge: 100%
Ending charge: 82%
Currently charging: yes
Opportunity rover used: yes
Hours: 121.9
Beginning charge: –
Ending charge: –
Currently charging: yes
Curiosity rover used: yes
Hours: 232.2
Beginning charge: 100%
Ending charge: 68%
Currently charging: yes
Perseverance rover used: no
Hours: 262.0
Beginning charge: –
Ending charge: –
Currently charging: yes
General notes on rovers: Duck tape on Curiosity’s seat is melting down and sticks to crew’s jumpsuits.
Summary of Hab operations:
WATER USE: 25.5 gal (from 6PM Sol9 to 6PM Sol10)
Water (static tank): 264.5 gal (at 6:00PM)
Static tank pipe heater: off
Static tank heater: off
Toilet tank emptied: yes
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: 1,7,8 and 10
Radios: nominal
All radio and spacesuit batteries checked before EVA: Yes. All nominal.
Summary of GreenHab operations:
WATER USE: 14 gallons (+1 gallon for Science Dome)
Heater: Off
Supplemental light: Off
Harvest: 105 gr salad mix sprouts (from crew)
Summary of ScienceDome operations: Laia Ribas, Neus Sabaté and Cesca Cufí have been working on their experiments. Neus Sabaté has used 50 ml of HCl (10%).
Dual split: Heat or AC, Off
Summary of RAM operations: Ariadna Farrés used a work bench to mount a sextant.
Summary of any observatory issues:
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues: none
Summary of other system’s issues: none
Questions, concerns and requests to Mission Support:
Questions: none
Kitchen: knobs replacement – expected this week

Journalist Report – April 26th

Crew 280 Journalist Report 26-04-2023

Author: Núria Jar, Crew Journalist

The crew members of Hypatia I were relieved to enjoy a warmer day on Mars as the temperature climbed above 20 degrees Celsius, following some chilly days at the beginning of their mission.

On average, the temperature on Mars is -63ºC, but the difference between Martian day and night goes from 35ºC during the day and drops significantly at night to -80ºC. The extreme temperature variation can make exploration and survival on Mars challenging. However, Hypatia I crew is surviving efficiently.

The Crew Scientist of Hypatia I, Ariadna Farrés Basiana, demonstrated her resourcefulness by making a homemade sextant using only a few objects to determine the latitude at the Mars Research Desert Station (MDRS). She took a rule, a piece of cord, one rock and a homemade protractor. The crew is looking forward to testing it out during the night to see how accurate the homemade sextant is in determining their location, in relation to the Polar Star.

Mars has an inclination that is very similar to Earth’s. So, the method used by the Crew Scientist to determine the latitude at the MDRS using a homemade sextant should work on Mars just as it does on Earth.

The extravehicular activity (EVA) of the day took place along Cow Dung Road 0110, just before reaching the Martian moons around Outlook. The long distance between this destination and the MDRS made radio communications between the crew and the base challenging. To address this issue, the Crew Commander, Mariona Badenas-Agustí, pointed the radio antenna towards the south to improve communication with the base. Fortunately, it worked and the EVA was a nice experience in the MDRS surroundings.

During the EVA, the GreenHab Officer, Laia Ribas, took the opportunity to collect sand samples from various locations. She is using microBIOMETER, a device that allows her to test the soil microbes on site, to study the possibility of life on Mars. The GreenHab Officer is leading this experiment in collaboration with the school of their kids, so the next generation is getting involved in scientific exploration and discovery. She finally discovered the presence of life on the Red Planet.

GreenHab Report – April 26th

Crew 280 Green Hab Report 26Apr2023

SOL 10 Greenhab Report

GreenHab Officer: Laia Ribas

Environmental control: Open during the day and heater during the night

Temperatures: 78.5 F

Hours of supplemental light: N/A

Daily water usage for crops: 14 gallons

Daily water usage for research and/or other purposes: 1 gallons

Water in Blue Tank: 64 gallons

Time(s) of watering for crops: 08.30, 18:15 h

Changes to crops: N/A

The general watering of plants and crops was done. The door was open all day due to high temperatures. The temperature reached 82.5F at 14.00h. The floor has been swept.

Harvest: 105 g of Salad Mix sprouts belonging to the project named as “Seed for the Crew” led by the Green Hab Officer was harvested. These seeds are growing in an independent box with wet tissue paper.

Support/supplies needed.

EVA Report – April 26th

Crew 280 EVA Report 26-04-2023

EVA #10

Author: Carla Conejo (Executive Officer & Biologist)

The EVA had the following goals:
Documenting the EVA for outreach purposes, by Núria Jar (Journalist).
Drone flight recording, by Carla Conejo (Executive Officer & Biologist) and the support of Mariona Badenas (Commander & Astronomer).
Collecting sand samples for the CatMart project, by Laia Ribas (GreenHab Officer).

Start time: 10:10 AM

End time: 12:50 PM


After performing 5 minutes of depressurization in the front airlock, the EVA team took Curiosity and Spirit rovers and headed north towards the cross with Brahe Highway 1572. The crew performing the EVA had the goals of: i) collecting sand samples searching for Martian biosignatures for Laia Ribas’ (GreenHab Officer) project; ii) documenting the latter process with a GoPro camera; and iii) recording some Martian landscape images with a DJI Mini 3 Pro drone.

The communications with the base were established with normality during the whole EVA with checks-ins every 30’ minutes. Just before arriving at Brahe Highway 1572, we parked the rovers and the EVA team climbed a hill nearby next to the road. From there, communications were established with the base to inform of our position. After that, to facilitate the communications between the EVA crew and the EVA leader, the EVA leader decided to switch between channel one and two on the radios. After taking a 5-minutes rest, the EVA team split themselves into two groups. The first group, Mariona Badenas (Commander & Astronomer) and Carla Conejo (Executive Officer & Biologist) stayed in the hill to fly the drone and take some zenith pictures of Mars; whereas the second group, Laia Ribas (GreenHab Officer) and Núria Jar (Journalist) descended the hill to access a dry water drain to search for Martian biosignatures. The two groups maintained constant visual and radio communications.

Around 12:10PM the team decided to return to the MDRS, which they reached at 12:50PM. Before entering the Hab, we followed the 5-minutes pressurization protocol.

Destination: Cow Dung Road 0110 cross with Brahe Highway 1572. Our initial plan was to reach Overlook, but we finally explored the Brahe Highway 1572 area by foot.

Coordinates (use UTM WSG 84): 516000 E, 4255000 N

EVA participants: 4: Mariona Badenas (Commander & Astronomer) (EVA Leader), Carla Conejo (Executive Officer & Biologist), Laia Ribas (GreenHab Officer) and Núria Jar (Journalist).

Road(s) and routes per MDRS Map: Drive east on Cow Dung Road 0110 and park before arriving at Brahe Highway 1572.

Mode of travel: Drive east on Cow Dung Road 0110 and to Brahe Highway 1572, to walk around the area no farther than 1000m from the rovers.

Copyright © The Mars Society. All rights reserved. | Main Site